Vapour pressure of volatile liquid a is 20 MM of hg equimolar solution containing a and other volatile liquid b in solution vapour pressure of this solution is 45 mm of HG now another solution A and B has a vapour pressure 22.5 mm of HG find ratio of mole fraction of A and B in final solution

Grade:12

1 Answers

Abhiraj Pathak
5 years ago
Pao = 20mmHgPbo = 45mmHgP = 22.5mmHgP = Pao*xa +Pbo*xbxa + xb = 1xa = 1-xb Substituting the values22.5 = 20(1-xb) + 45xb on solvingxb = 0.1; xa = 0.9Thus xa/xb = 9